OpenStudy (anonymous):

You are playing Monopoly with a friend. She owns and has developed Park Place and Boardwalk with hotels. If you roll an 8, you will land on Park Place and owe her $1500. You will be able to pay, but you will not have money left afterward. If you roll a 9, you will land on Luxury Tax and have to pay $75. If you roll a 10, you will land on Boardwalk and owe her $2000, which is more than you can pay—and lose the game. If you roll a 7 or less, you will not have to pay any money to anyone. Your friend offers you insurance. Pay her $500 before you roll and even if you land on Boardwalk or Park

OpenStudy (anonymous):

i guess the question is should you do it

OpenStudy (anonymous):

not sure how to asses the value of losing the game so it is an open ended question your expected loss is the probability you lose times the amount you lose \[\frac{5}{36}\times 1500+\frac{7}{36}\times 75+\frac{3}{36}\times 2000\]but the last one is confusing because you will lose if you roll it and the game will be over